Planning week 21-28 July 2012

At the start of the week the flare-productive active region complex AR11519,520,521 had just passed around the west limb, leaving the quiescent active region AR 11524/25 close to the central meridian, and the also-quiet AR 11526 on the east limb. Later in the week more active regions appeared around the east limb, in particular AR 11527 and 528 in the north, and AR 11529 and 530 in the south. The default active region pointing changed from AR 11525 in OP1, AR 11526 in OP2 to AR 11530 in OP3.

HOP 81 was supported on 22nd and 23rd. HOP 173 was run every day from 22nd onwards but EIS support was not requested and I did not give it (the observing period was awkward, with two 2 hour chunks with an XRT synoptic in between). HOP 79 was scheduled for the 26th with a north-south scan.

Since I'm the CO, the first runs of the EIS-COMP joint study were scheduled for the 21st and 24th, the east limb being the target in both cases.



Saturday 21 July (OP 1)

AIA 193 image

My plan for this OP was to monitor AR 11525 with the large AR study GDZ_300x384_S2S3_35s every 7 hours or so. 10 rasters were scheduled, although one was swapped for an Atlas_60 study.

The EIS-COMP study was run at the east limb at 18:43 (north) and 20:09 (south).

Sunday 22 July (OP1)

AIA 193 image

Monitoring of AR 11525 continued.

HOP 81 was supported for 10:45-15:35, with a north limb pointing.

Monday 23 July (OP 1)

AIA 193 image

Monitoring of AR 11525 continued.

HOP 81 was supported for 09:43-14:33, with an east limb pointing.

For the period 18:28-22:23, a flare study was performed on AR 11525 in the vain hope that there may be some small flares (it doesn't look like there was).

Tuesday 24 July (OP 1)

AIA 193 image

Monitoring of AR 11525 continued.

Tuesday 24 July (OP 2)

AIA 193 image

Pointing switched to AR 11526, which crosses the central meridian during this OP. I continued scheduling GDZ_300x384_S2S3_35s every 7 hours or so, with 7 rasters in all.

The COMP study was run again at 17:23 (north) and 18:49 (south). Note there are two ARs at the limb for this observation, with some interconnecting loops so this might be a good data-set.

Wednesday 25 July (OP 2)

AIA 193 image

Monitoring of AR 11526 continued.

The north pointing of the Hinode co-alignment program was performed at 22:27.

Thursday 26 July (OP 2)

Monitoring of AR 11526 continued.

The east pointing of the Hinode co-alignment program was performed at 00:15.

There are quite a lot of nice footpoints in AR 11526, so I scheduled the footpoint hunter study at 05:01, with the response being PRY_footpoints_HI2. I'm bit concerned how much of this data will come back (assuming the trigger goes off). Note that I scheduled this study for when the AR was close to the meridian.

Thursday 26 July (OP 3)

The north-south scan of HOP 79 was performed from 09:55 to 02:40 on 27 July.

Friday 27 July (OP 3)

AIA 193 image

HOP 79 support continued until 02:40.

AR 11530 was monitored with four runs of the GDZ_300x384_S2S3_35s study scheduled at 05:38, 13:51, 20:24, and 06:15 (28 July).

The Cam_flare_diag study was run during four consecutive day time orbits from 07:20 to 13:23. It was difficult to select an "active" part of the AR, so I just chose the region brightest in AIA 335.

A second run of Cam_flare_diag was scheduled for three further consecutive orbits from 22:06 to 02:31 (28 July).

Saturday 28 July (OP 3)

Monitoring of AR 11530 and the Cam_flare_diag study were continued.